OpenStudy (anonymous):

what is a rule for the function represented by (0, 2/19), (1, 1 2/19), (2, 4 2/19), (3, 9 2/19), (4, 16 2/19), and (5, 25 2/19)? Explain your reasoning

OpenStudy (anonymous):

Notice \[ f(0)=\frac 2 {19} = 0^2 +\frac 2 {19}\\ f(1)=1+ \frac 2 {19} = 1^1 + \frac 2 {19}\\ f(2)=4+ \frac 2 {19} = 2^2 + \frac 2 {19}\\ f(3)=9+ \frac 2 {19} = 3^2 + \frac 2 {19} \] What can you conclude?

OpenStudy (anonymous):

\[ f(4) = 16 + \frac 2 {19} = 4^2 + \frac 2 {19}\\ f(5) = 25 + \frac 2 {19} = 5^2+\frac 2 {19} \]

OpenStudy (anonymous):

So \[ f(x)= x^{?} + ? \]